# A Survey of Safe Reinforcement Learning and Constrained MDPs: A Technical Survey on Single-Agent and Multi-Agent Safety

### v1

@@ Abstract
Safe Reinforcement Learning (SafeRL) is the subfield of reinforcement learning that explicitly deals with safety constraints during the learning and deployment of agents. This survey provides a mathematically rigorous overview of SafeRL formulations based on Constrained Markov Decision Processes (CMDPs) and extensions to Multi-Agent Safe RL (SafeMARL). We review theoretical foundations of CMDPs, covering definitions, constrained optimization techniques, and fundamental theorems. We then summarize state-of-the-art algorithms in SafeRL for single agents, including policy gradient methods with safety guarantees and safe exploration strategies, as well as recent advances in SafeMARL for cooperative and competitive settings. Additionally, we propose five open research problems to advance the field, with three focusing on SafeMARL. Each problem is described with motivation, key challenges, and related prior work. This survey is intended as a technical guide for researchers interested in SafeRL and SafeMARL, highlighting key concepts, methods, and open future research directions.
@@ 1 Introduction
Definition 1.1.
The goal in SafeRL is to maximize performance (cumulative reward) while satisfying safety constraints during training and deployment.
A common framework for SafeRL is the Constrained Markov Decision Process (CMDP) introduced by Altman (1999). In a CMDP, an agent seeks to maximize expected return subject to one or more constraints (e.g., bounds on certain costs or probabilities of unsafe events). This framework allows formalizing safety requirements as mathematical constraints and provides tools from constrained optimization and control theory to enforce them. SafeRL algorithms often leverage CMDP theory to find policies that respect constraints (at least approximately) while learning efficiently. SafeRL has gained significant attention in recent years. Early work in SafeRL explored modifications of the RL objective to encode risk or safety (e.g., worst-case guarantees Heger (1994), risk-sensitive criteria Borkar (2002), or probability of failure constraints Geibel and Wysotzki (2005)). More recent approaches explicitly enforce constraints during learning using techniques like Lagrange multipliers, trust-region methods, or safety monitors. There have been comprehensive surveys of SafeRL (e.g., Garcia and Fernandez (2015)) and increasing theoretical study of constrained RL algorithms Achiam et al. (2017); Chow et al. (2018a). An emerging frontier is Multi-Agent Safe Reinforcement Learning (SafeMARL), which considers multiple agents learning and interacting under safety constraints. SafeMARL is crucial for applications like coordinated robotics, drone swarms, and autonomous driving with multiple vehicles, where safety conditions involve interactions among agents. SafeMARL introduces additional challenges such as coordinating safety in a team, handling the coupling of constraints across agents, and new solution concepts (like safe equilibria Ganzfried (2022) in competitive settings). While single-agent SafeRL is relatively well-studied, SafeMARL remains a young research area with many open problems ElSayed-Aly et al. (2021); Gu et al. (2023). This survey provides
@@ 2 Related Work
Our survey focuses primarily on the constrained criterion approach (CMDPs), which has become the prevalent formalism for SafeRL in recent years. Since 2015, the field has advanced with new algorithms and theoretical results. Recent reviews such as Wachi et al. Wachi et al. (2024) examine various formulations of safety constraints (e.g., how constraints are represented and enforced) and draw connections between them. Another forthcoming survey by Gu et al. Gu et al. (2024) provides an extensive review of SafeRL methods, theory, and applications, reflecting the growing maturity of the field. These works indicate an increasing interest in unifying SafeRL concepts and developing a systematic understanding of safety constraint representations and their implications. On the multi-agent side, SafeMARL has been less surveyed due to its emergent status. Gu et al. Gu et al. (2023) investigate safe multi-robot control tasks and propose algorithms like Multi-Agent Constrained Policy Optimization (MACPO). Some recent papers introduce safe multi-agent learning algorithms or frameworks ElSayed-Aly et al. (2021); Gu et al. (2023); Zhang et al. (2024), but a comprehensive survey of SafeMARL is still lacking. Our work contributes by reviewing both single-agent and multi-agent safe RL in one document and highlighting SafeMARL-specific challenges. Other related areas include robust RL (handling model uncertainty or adversarial disturbances) and reward hacking / alignment (ensuring the specified reward leads to intended safe behavior). While robust RL (e.g., solving worst-case MDPs) and SafeRL share some techniques (like min-max optimization), they address different problem formulations (uncertainty vs. explicit constraints). Similarly, reward specification and alignment problems are complementary to SafeRL: one can combine learned reward shaping with SafeRL constraints to yield agents that both seek correct objectives and stay safe Amodei et al. (2016); Achiam et al. (2017). We touch upon these connections where relevant. In summary, our survey builds upon and extends prior work by providing a focused treatment of CMDP-based SafeRL and the novel area of SafeMARL, presented in a rigorous yet accessible manner for researchers.
@@ 3.1 Markov Decision Processes (MDPs)
We begin with the standard Markov Decision Process (MDP) formulation of an RL problem. An MDP is defined by the tuple , where
is a (finite or continuous) set of states.
is a set of actions available to the agent.
is the transition probability function (Markovian dynamics), giving the distribution over next states when action is taken in state .
is a reward function (or including next state, depending on context) giving a scalar reward for executing action in state .
is a discount factor that weights immediate vs. future rewards (with typically for infinite-horizon problems).
A (stationary) policy is a mapping from states to a distribution over actions. We denote as the probability of taking action in state under .
The value function for a policy is
the expected cumulative discounted reward starting from state and following . The goal in standard RL is to find an optimal policy maximizing for all (or maximizing a specific initial state or distribution performance). Equivalently, one maximizes the return for some initial state distribution . In unconstrained RL, solves .
@@ 3.2 Constrained Markov Decision Processes (CMDPs)
is called the feasible policy set. We assume this set is non-empty (the constraints are attainable). Problem (1) is the standard formulation of SafeRL as a CMDP optimization problem Altman (1999). It is a constrained Markov decision problem which, in principle, can be solved via dynamic programming or linear programming if the model is known and state-action spaces are small. Eitan Altman’s foundational work Altman (1999) established that for finite CMDPs, there exists an optimal policy that is stationary (time-independent) and, if multiple constraints are present, possibly stochastic (randomized). Intuitively, sometimes a mixture of actions is required to exactly satisfy multiple constraints: a deterministic policy might violate a constraint, whereas a stochastic policy can blend strategies to meet the constraint bounds exactly.
@@
which are also linear in . This linear structure is crucial as it allows the CMDP optimization problem to be expressed as a linear program (LP).
The CMDP can then be written as a linear program:
@@ 3.3 Constraint Types and Safety Specifications
Robotics — Torque or Force Limits: Robotic manipulators and mobile robots have strict actuator limits. A common constraint is , where is the torque vector applied at time . Exceeding these limits even once can cause irreversible hardware damage. Therefore, the torque constraint must hold at every step.
@@
Autonomous Driving — Collision Avoidance: Autonomous vehicles must avoid collisions at all times. This is often modeled as a minimum distance constraint, , where is a safety margin. Unlike reward penalties for collisions, instantaneous constraints aim to ensure that no collision ever occurs, even during learning.
Aerial Vehicles (Drones) — Altitude Constraints: Drones often operate within restricted altitude corridors, leading to constraints of the form . Exceeding altitude boundaries may result in collisions with terrain (if ) or violation of airspace regulations (if ). Such constraints must be enforced at all times.
@@
Medical Applications — Dose Limits in Treatment Planning: In adaptive radiation therapy or drug administration, instantaneous dosage constraints are essential. The instantaneous constraint may take the form , limiting the maximum dose administered at each step to prevent severe side effects.
@@
Power Systems — Voltage and Current Limits: Power grids are subject to operational safety limits such as for voltage levels or similar constraints on current. Violations could cause system instability, equipment damage, or even large-scale blackouts. Safe control must respect these constraints instantaneously.
@@
Industrial Process Control — Pressure Limits: In chemical plants, nuclear reactors, and manufacturing systems, pressure constraints of the form are typical. Exceeding pressure thresholds even once may lead to catastrophic failures such as explosions or hazardous material leaks.
@@
Spacecraft and Autonomous Vehicles — Safe Landing or Docking Probability Blackmore et al. (2010); Ono and Williams (2015, 2013); Chow et al. (2015): In space missions or autonomous landing scenarios, failure is often defined as crashing during landing or docking. One may enforce a constraint such as , where is a small acceptable risk level. The cost function is defined as if leads to a crash state.
@@
Specifically, The paper Blackmore et al. (2010) is one of the earliest works in chance-constrained motion planning and is frequently cited in spacecraft and UAV planning and Ono and Williams (2015), is a classic reference on chance-constrained formulations for spacecraft landing and docking. The paper Ono and Williams (2013) directly deals with probability of failure constraints for spacecraft control.
@@
Robotics — Falling or Tipping Over Berkenkamp et al. (2017); Wabersich and Zeilinger (2018); Berkenkamp and Schoellig (2015); Thananjeyan and et al. (2021): In humanoid or legged robots, failure is typically associated with falling down. The agent is required to maintain to ensure physical integrity and task feasibility. In this case, any state classified as “fallen” is marked as a failure state, and if the next state is a fallen state. Specifically, Berkenkamp et al. (2017) is a classic paper that specifically addresses falling in legged robots and balance maintenance as a safety constraint. They model unsafe states (like falls) and ensure with high probability that they are avoided. The paper Wabersich and Zeilinger (2018) introduces a safety certification approach ensuring that robots do not enter dangerous states (including falls). It applies to both wheeled and legged robots. The earlier work Berkenkamp and Schoellig (2015) focuses on safe policy learning for balancing and preventing falls. It explicitly models unsafe states (falls) in the dynamics and safety set. Thananjeyan and et al. (2021) addresses falling probability directly in RL-based locomotion. It models type constraints explicitly for quadruped robots.
@@
Healthcare — Patient Mortality or Critical Failure: In reinforcement learning for clinical decision-making (e.g., ICU treatment policies), a failure may be defined as the patient’s mortality or reaching a critical medical condition. The constraint limits the treatment policy to maintain acceptable risk levels. Here, failure states correspond to medical emergencies.
@@
In particular, Raghu et al. (2017), models ICU treatment as an MDP where mortality is treated as an absorbing failure state. While optimizing expected return, they explicitly consider trajectories leading to death. In 2018, Komorowski et al. (2018) proposed one of the first full-scale RL systems for ICU treatment. It models mortality as an outcome, and the trained policies aim to reduce the occurrence of death. Following this in Gottesman et al. (2019), proposed a foundational paper outlining safety and interpretability concerns in clinical RL. It explicitly discusses mortality and adverse outcomes as critical failure events. While not formalizing constraints as it motivates their necessity. In the same year, Jiang et al. (2019) explored the importance of constraint enforcement for safe clinical decision-making. It explicitly addresses risk of critical events in clinical policies.
@@
Finance — Bankruptcy or Insolvency Events Jiang et al. (2019); Filar et al. (1995); Borkar and Jain (2014); Chow et al. (2015); Huang et al. (2018); Ravanbakhsh et al. (2019): In financial portfolio management, the failure event could be the agent’s wealth dropping below a bankruptcy threshold. The probability of this event is often constrained by to limit risk exposure. The cost function is if wealth crosses the bankruptcy boundary.
@@
The paper Filar et al. (1995) is one of the early papers introducing risk-sensitive portfolio optimization with Markov decision processes. It addresses ruin (bankruptcy) probabilities explicitly. In 2014 paper Borkar and Jain (2014), proposes to directly deals with probability of ruin (bankruptcy) in constrained MDPs. It proposes algorithms under constraints like The paper Huang et al. (2018) models percentile-based risk for financial RL tasks where falling below a wealth threshold triggers bankruptcy. While applied to cloud scheduling, Ravanbakhsh et al. (2019) demonstrates the same probability of ruin modeling, similar to financial insolvency constraints.
@@
Manufacturing — Production System Breakdown: In industrial automation, a failure might occur when production machinery exceeds thermal, mechanical, or chemical safety limits leading to breakdown. The probability of such system failure is constrained to be below a pre-specified threshold, e.g., .
@@
Power Grids — Blackout Events: In power system control, blackouts (large-scale power failures) are often modeled as absorbing failure states. The system may enforce to reduce the chance of a cascading failure. Failure is usually caused by overloading, component failures, or instability.
@@
Autonomous Driving — Variance-Constrained Driving Comfort: While avoiding collisions is a safety constraint, maintaining comfortable driving also involves controlling the variance of acceleration, jerk, or lane deviations. A variance constraint of the form can ensure that passenger discomfort due to aggressive or unstable maneuvers remains limited, reducing the risk of loss of control or accidents.
Specifically, in 2012, Tamar et al. (2012), introduced variance-constrained reinforcement learning where variance of return is explicitly controlled. It is applicable to driving scenarios when controlling variance of acceleration or jerk. Zhu et al. (2020) explicitly focuses on reducing control variability to improve smoothness and driving comfort. In 2021 Kiran et al. (2021) wrote a comprehensive survey that discusses driving comfort (acceleration, jerk minimization, smoothness) as key secondary objective in autonomous driving, and mentions various papers that incorporates constraints.
Finance — CVaR-Constrained Portfolio Optimization: In financial portfolio management, it is common to limit the Conditional Value-at-Risk (CVaR) of the portfolio’s return. A CVaR constraint ensures that the expected loss in the worst of cases does not exceed a tolerable threshold. This is widely used to manage downside risk beyond what variance alone captures.
@@
Healthcare — CVaR for Adverse Outcomes: In healthcare applications such as treatment planning or resource allocation, minimizing the expected number of adverse events may not be sufficient. A CVaR constraint on cumulative adverse events or side effects ensures that treatment policies control the likelihood of rare but severe negative outcomes.
Although general, the paper Prashanth and Ghavamzadeh (2016) is frequently cited in healthcare RL as it provides risk-sensitive methods including CVaR for controlling adverse events. Mattei and Gopalan (2019) specifically applies CVaR-based reinforcement learning to the problem of sepsis treatment, modeling adverse outcomes such as mortality and organ failure. This is one of the most direct applications of CVaR in healthcare. A position paper Gottesman et al. (2019) emphasizes that minimizing expected adverse outcomes is insufficient and recommended the use of risk measures (e.g., CVaR). Although it doesn’t present an algorithm, it motivates CVaR as a necessary tool in treatment planning.
@@
Supply Chain Management — Risk-Averse Inventory Control: In supply chains, stockout events (inventory drops below zero) cause disruptions. Instead of just minimizing expected stockouts, a CVaR constraint on stockout penalties ensures that even in rare demand spikes, the risk of large cumulative stockouts is controlled.
@@
Robotics — CVaR-Constrained Trajectory Optimization: For autonomous robots navigating uncertain environments, one may use CVaR constraints on cumulative collision risk or energy consumption. This ensures that the robot does not just minimize average risk but is also robust against worst-case environmental uncertainties or adversarial perturbations.
The paper Majumdar and Pavone (2020) directly studies CVaR-constrained trajectory optimization for robots under uncertainty. Formulates trajectory optimization problems where collision risk is controlled using CVaR. The paper Zhang et al. (2020) focuses on risk-averse path planning under environmental uncertainty using CVaR. Provides algorithms and examples for safe robot navigation with collision risk control. The paper Singh et al. (2018) considers autonomous driving (viewed as mobile robotics) with risk-sensitive trajectory optimization. It uses CVaR to constrain collision probabilities and improve robustness to adversarial perturbations.
Power Systems — Risk-Sensitive Stability Control: In power grid operations, rather than just minimizing expected frequency deviations or power outages, operators may use CVaR or variance constraints to ensure that the probability of large-scale instabilities remains acceptably low, accounting for rare but impactful demand or supply fluctuations.
The paper Bitar and Low (2012) addresses reliability and demand uncertainties in power systems with a risk-sensitive approach. It Models constraints on load shedding and supply-demand balancing. The paper Dall’Anese et al. (2015) directly introduces chance-constrained optimization for voltage stability and power flow, which is equivalent to controlling the probability of instability; CVaR and probability bounds are discussed. In 2012, Jiang et al. (2012) introduces risk-constrained OPF formulations using CVaR and chance constraints. It ensures that the probability of voltage violations and instabilities is below a prescribed risk level. A year earlier in 2011, Zhang et al. (2011) models the variance and tail risk of power system instability due to fluctuating wind generation. While not CVaR directly, it motivates variance and higher-order moment-based risk constraints.
@@
Multi-objective viewpoint: SafeRL can be seen as a multi-objective optimization where one objective is reward and others are (negative) costs. The constraint formulation picks one point on the Pareto frontier by treating costs as hard constraints. Alternatively, one could combine reward and costs into a single scalar reward via weighted sum (penalty method), but that requires tuning weights and does not guarantee constraint satisfaction Achiam et al. (2017). Constrained formulation cleanly separates objectives and safety.

### v2

@@ 1 Introduction
A common framework for SafeRL is the Constrained Markov Decision Process (CMDP) introduced by Altman (1999). In a CMDP, an agent seeks to maximize expected return subject to one or more constraints (e.g., bounds on certain costs or probabilities of unsafe events). This framework allows formalizing safety requirements as mathematical constraints and provides tools from constrained optimization and control theory to enforce them. SafeRL algorithms often leverage CMDP theory to find policies that respect constraints (at least approximately) while learning efficiently. SafeRL has gained significant attention in recent years. Early work in SafeRL explored modifications of the RL objective to encode risk or safety (e.g., worst-case guarantees Heger (1994), risk-sensitive criteria Borkar (2002), or probability of failure constraints Geibel and Wysotzki (2005)). More recent approaches explicitly enforce constraints during learning using techniques like Lagrange multipliers, trust-region methods, or safety monitors. There have been comprehensive surveys of SafeRL (e.g., Garcia and Fernandez (2015)) and increasing theoretical study of constrained RL algorithms Achiam et al. (2017); Chow et al. (2018). An emerging frontier is Multi-Agent Safe Reinforcement Learning (SafeMARL), which considers multiple agents learning and interacting under safety constraints. SafeMARL is crucial for applications like coordinated robotics, drone swarms, and autonomous driving with multiple vehicles, where safety conditions involve interactions among agents. SafeMARL introduces additional challenges such as coordinating safety in a team, handling the coupling of constraints across agents, and new solution concepts (like safe equilibria Ganzfried (2023) in competitive settings). While single-agent SafeRL is relatively well-studied, SafeMARL remains a young research area with many open problems ElSayed-Aly et al. (2021); Gu et al. (2023). This survey provides
@@ 2 Related Work
Our survey focuses primarily on the constrained criterion approach (CMDPs), which has become the prevalent formalism for SafeRL in recent years. Since 2015, the field has advanced with new algorithms and theoretical results. Recent reviews such as Wachi et al. Wachi et al. (2024) examine various formulations of safety constraints (e.g., how constraints are represented and enforced) and draw connections between them. Another forthcoming survey by Gu et al. Gu et al. (2024) provides an extensive review of SafeRL methods, theory, and applications, reflecting the growing maturity of the field. Domain-specific surveys have also emerged, such as Brunke et al. Brunke et al. (2022) who provide a comprehensive review of safe RL methods for robotics. These works indicate an increasing interest in unifying SafeRL concepts and developing a systematic understanding of safety constraint representations and their implications. On the multi-agent side, SafeMARL has been less surveyed due to its emergent status. Gu et al. Gu et al. (2023) investigate safe multi-robot control tasks and propose algorithms like Multi-Agent Constrained Policy Optimization (MACPO). Some recent papers introduce safe multi-agent learning algorithms or frameworks ElSayed-Aly et al. (2021); Gu et al. (2023); Zhang et al. (2024), but a comprehensive survey of SafeMARL is still lacking. Our work contributes by reviewing both single-agent and multi-agent safe RL in one document and highlighting SafeMARL-specific challenges. Other related areas include robust RL (handling model uncertainty or adversarial disturbances) and reward hacking / alignment (ensuring the specified reward leads to intended safe behavior). While robust RL (e.g., solving worst-case MDPs) and SafeRL share some techniques (like min-max optimization), they address different problem formulations (uncertainty vs. explicit constraints). Similarly, reward specification and alignment problems are complementary to SafeRL: one can combine learned reward shaping with SafeRL constraints to yield agents that both seek correct objectives and stay safe Amodei et al. (2016); Achiam et al. (2017). Benchmark suites such as the AI Safety Gridworlds Leike et al. (2017) and SafeLife Wainwright and Eckersley (2021) specifically test for specification robustness and side-effect avoidance. We touch upon these connections where relevant. In summary, our survey builds upon and extends prior work by providing a focused treatment of CMDP-based SafeRL and the novel area of SafeMARL, presented in a rigorous yet accessible manner for researchers.
@@ 3.2 Constrained Markov Decision Processes (CMDPs)
is called the feasible policy set.
We assume this set is non-empty (the constraints are attainable). Problem (1) is the standard formulation of SafeRL as a CMDP optimization problem Altman (1999). It is a constrained Markov decision problem which, in principle, can be solved via dynamic programming or linear programming if the model is known and state-action spaces are small. Eitan Altman’s foundational work Altman (1999) established that for finite CMDPs, there exists an optimal policy that is stationary (time-independent) and, if multiple constraints are present, possibly stochastic (randomized). Intuitively, sometimes a mixture of actions is required to exactly satisfy multiple constraints: a deterministic policy might violate a constraint, whereas a stochastic policy can blend strategies to meet the constraint bounds exactly.
@@
which are also linear in . This linear structure is crucial as it allows the CMDP optimization problem to be expressed as a linear program (LP). The CMDP can then be written as a linear program:
@@ 3.3 Constraint Types and Safety Specifications
Robotics –Torque or Force Limits Dalal et al. (2018); Cheng et al. (2019); Achiam et al. (2017): Robotic manipulators and mobile robots have strict actuator limits. A common constraint is , where is the torque vector applied at time . Exceeding these limits even once can cause irreversible hardware damage. Therefore, the torque constraint must hold at every step. Dalal et al. Dalal et al. (2018) proposed a safety layer that projects RL actions to satisfy such constraints, while Cheng et al. Cheng et al. (2019) combined model-free RL with control barrier functions to enforce actuator limits during learning.
@@
Autonomous Driving –Collision Avoidance Shalev-Shwartz et al. (2018); Isele et al. (2018); Nguyen and Han (2023): Autonomous vehicles must avoid collisions at all times. This is often modeled as a minimum distance constraint, , where is a safety margin. Unlike reward penalties for collisions, instantaneous constraints aim to ensure that no collision ever occurs, even during learning. Shalev-Shwartz et al. Shalev-Shwartz et al. (2018) formalized hard safety constraints via the Responsibility-Sensitive Safety (RSS) framework, while Isele et al. Isele et al. (2018) used prediction-based constraints to safely learn intersection-handling behaviors.
Figure 4: SafeRL for autonomous driving. The RL agent receives state observations from sensors and outputs actions (steering, acceleration). A safety filter ensures the executed action satisfies the instantaneous constraint , preventing collisions while optimizing travel efficiency. Aerial Vehicles (Drones) –Altitude Constraints Fisac et al. (2019); Gillula and Tomlin (2012); Yuan et al. (2022): Drones often operate within restricted altitude corridors, leading to constraints of the form . Exceeding altitude boundaries may result in collisions with terrain (if ) or violation of airspace regulations (if ). Such constraints must be enforced at all times. Fisac et al. Fisac et al. (2019) proposed a Hamilton-Jacobi reachability-based framework guaranteeing state constraint satisfaction for quadrotors, and Gillula and Tomlin Gillula and Tomlin (2012) demonstrated guaranteed safe online learning on a quadrotor with altitude bounds.
@@
Medical Applications –Dose Limits in Treatment Planning Tseng et al. (2017); Sprouts et al. (2022): In adaptive radiation therapy or drug administration, instantaneous dosage constraints are essential. The instantaneous constraint may take the form , limiting the maximum dose administered at each step to prevent severe side effects. Tseng et al. Tseng et al. (2017) developed a deep RL framework for dose fractionation in lung cancer constrained by tissue complication limits, and Sprouts et al. Sprouts et al. (2022) trained a DRL-based treatment planner enforcing hard dose-volume constraints on organs at risk.
@@
Figure 5: SafeRL for healthcare treatment planning. The RL policy observes patient state (vitals, lab values) and prescribes treatment actions (drug dosage, ventilator settings). A safety constraint enforces dose limits () and bounds mortality risk (). Unsafe actions are rejected and the patient state evolves based on the administered treatment. Power Systems –Voltage and Current Limits Vu et al. (2021); Duan et al. (2020): Power grids are subject to operational safety limits such as for voltage levels or similar constraints on current. Violations could cause system instability, equipment damage, or even large-scale blackouts. Safe control must respect these constraints instantaneously. Vu et al. Vu et al. (2021) proposed barrier function-based safe RL for emergency voltage control with hard safety bounds, while Duan et al. Duan et al. (2020) developed a DRL-based autonomous voltage control agent that maintains voltage within operational limits.
@@
Industrial Process Control –Pressure Limits Kim and Kim (2022): In chemical plants, nuclear reactors, and manufacturing systems, pressure constraints of the form are typical. Exceeding pressure thresholds even once may lead to catastrophic failures such as explosions or hazardous material leaks. Kim and Oh Kim and Kim (2022) developed safe model-based RL using Lyapunov barrier functions for chemical process control (CSTR) with hard state and input constraints on temperature and pressure.
@@
Spacecraft and Autonomous Vehicles –Safe Landing or Docking Probability Blackmore et al. (2006); Ono et al. (2015); Chow et al. (2015): In space missions or autonomous landing scenarios, failure is often defined as crashing during landing or docking. One may enforce a constraint such as , where is a small acceptable risk level. The cost function is defined as if leads to a crash state.
@@
Specifically, The paper Blackmore et al. (2006) is one of the earliest works in chance-constrained motion planning and is frequently cited in spacecraft and UAV planning and Ono et al. (2015), is a classic reference on chance-constrained formulations for spacecraft landing and docking. The paper Ono et al. (2015) directly deals with probability of failure constraints for spacecraft control.
@@
Robotics: Falling or Tipping Over Berkenkamp et al. (2017); Wabersich and Zeilinger (2018); Berkenkamp and Schoellig (2015): In humanoid or legged robots, failure is typically associated with falling down. The agent is required to maintain to ensure physical integrity and task feasibility. In this case, any state classified as “fallen” is marked as a failure state, and if the next state is a fallen state. Specifically, Berkenkamp et al. (2017) is a classic paper that specifically addresses falling in legged robots and balance maintenance as a safety constraint. They model unsafe states (like falls) and ensure with high probability that they are avoided. The paper Wabersich and Zeilinger (2018) introduces a safety certification approach ensuring that robots do not enter dangerous states (including falls). It applies to both wheeled and legged robots. The earlier work Berkenkamp and Schoellig (2015) focuses on safe policy learning for balancing and preventing falls. It explicitly models unsafe states (falls) in the dynamics and safety set.
@@
Healthcare: Patient Mortality or Critical Failure Raghu et al. (2018); Gottesman et al. (2019); Jia et al. (2020); Tu et al. (2025): In reinforcement learning for clinical decision-making (e.g., ICU treatment policies), a failure may be defined as the patient’s mortality or reaching a critical medical condition. The constraint limits the treatment policy to maintain acceptable risk levels. Here, failure states correspond to medical emergencies.
@@
In particular, Raghu et al. (2018), models ICU treatment as an MDP where mortality is treated as an absorbing failure state. While optimizing expected return, they explicitly consider trajectories leading to death. In 2018, M et al. (2024) did a comprehensive study on use of RL systems for ICU treatment. Gottesman et al. (2019), proposed a foundational paper outlining safety and interpretability concerns in clinical RL. It explicitly discusses mortality and adverse outcomes as critical failure events. While not formalizing constraints as it motivates their necessity.
@@
Finance: Bankruptcy or Insolvency Events Neto et al. (2020); Borkar and Jain (2014); Chow et al. (2015, 2017); Schlosser (2020): In financial portfolio management, the failure event could be the agent’s wealth dropping below a bankruptcy threshold. The probability of this event is often constrained by to limit risk exposure. The cost function is if wealth crosses the bankruptcy boundary.
@@
The paper Neto et al. (2020) discusses risk-sensitive portfolio optimization with Markov decision processes. It addresses ruin (bankruptcy) probabilities explicitly. In 2014 paper Borkar and Jain (2014), proposes to directly deals with probability of ruin (bankruptcy) in constrained MDPs. It proposes algorithms under constraints like The paper Chow et al. (2017) models percentile-based risk for financial RL tasks where falling below a wealth threshold triggers bankruptcy. While applied to cloud scheduling, demonstrates the same probability of ruin modeling, similar to financial insolvency constraints.
@@
Manufacturing –Production System Breakdown: In industrial automation, a failure might occur when production machinery exceeds thermal, mechanical, or chemical safety limits leading to breakdown. The probability of such system failure is constrained to be below a pre-specified threshold, e.g., .
@@
Power Grids –Blackout Events: In power system control, blackouts (large-scale power failures) are often modeled as absorbing failure states. The system may enforce to reduce the chance of a cascading failure. Failure is usually caused by overloading, component failures, or instability.
@@
Autonomous Driving –Variance-Constrained Driving Comfort: While avoiding collisions is a safety constraint, maintaining comfortable driving also involves controlling the variance of acceleration, jerk, or lane deviations. A variance constraint of the form can ensure that passenger discomfort due to aggressive or unstable maneuvers remains limited, reducing the risk of loss of control or accidents.
Specifically, in 2012, Tamar et al. (2012), introduced variance-constrained reinforcement learning where variance of return is explicitly controlled. It is applicable to driving scenarios when controlling variance of acceleration or jerk. Huang et al. (2026) explicitly focuses on reducing control variability to improve smoothness and driving comfort. In 2021 Kiran et al. (2022) wrote a comprehensive survey that discusses driving comfort (acceleration, jerk minimization, smoothness) as key secondary objective in autonomous driving, and mentions various papers that incorporates constraints.
Finance: CVaR-Constrained Portfolio Optimization: In financial portfolio management, it is common to limit the Conditional Value-at-Risk (CVaR) of the portfolio’s return. A CVaR constraint ensures that the expected loss in the worst of cases does not exceed a tolerable threshold. This is widely used to manage downside risk beyond what variance alone captures.
@@
Figure 6: SafeRL for risk-constrained portfolio management. The RL agent observes market state and outputs portfolio allocations. A CVaR constraint limits tail risk, ensuring the expected loss in the worst of scenarios stays bounded. The inset shows the loss distribution with the CVaR tail region highlighted. Healthcare: CVaR for Adverse Outcomes: In healthcare applications such as treatment planning or resource allocation, minimizing the expected number of adverse events may not be sufficient. A CVaR constraint on cumulative adverse events or side effects ensures that treatment policies control the likelihood of rare but severe negative outcomes.
Although general, the paper Prashanth and Ghavamzadeh (2013) is frequently cited in healthcare RL as it provides risk-sensitive methods including CVaR for controlling adverse events. A position paper Gottesman et al. (2019) emphasizes that minimizing expected adverse outcomes is insufficient and recommended the use of risk measures (e.g., CVaR). Although it doesn’t present an algorithm, it motivates CVaR as a necessary tool in treatment planning.
@@
Supply Chain Management –Risk-Averse Inventory Control: In supply chains, stockout events (inventory drops below zero) cause disruptions. Instead of just minimizing expected stockouts, a CVaR constraint on stockout penalties ensures that even in rare demand spikes, the risk of large cumulative stockouts is controlled.
@@
Robotics –CVaR-Constrained Trajectory Optimization: For autonomous robots navigating uncertain environments, one may use CVaR constraints on cumulative collision risk or energy consumption. This ensures that the robot does not just minimize average risk but is also robust against worst-case environmental uncertainties or adversarial perturbations.
The paper Ahmadi et al. (2022) directly studies CVaR-constrained trajectory optimization for robots under uncertainty (see also Hakobyan et al. (2019); Ahmadi et al. (2021); Bian et al. (2023) for CVaR-based motion planning and path planning). Formulates trajectory optimization problems where collision risk is controlled using CVaR. The paper Yu et al. (2026) focuses on risk-averse path planning under environmental uncertainty using CVaR. Provides algorithms and examples for safe robot navigation with collision risk control.
Power Systems –Risk-Sensitive Stability Control: In power grid operations, rather than just minimizing expected frequency deviations or power outages, operators may use CVaR or variance constraints to ensure that the probability of large-scale instabilities remains acceptably low, accounting for rare but impactful demand or supply fluctuations.
The paper Bitar and Xu (2017) addresses reliability and demand uncertainties in power systems with a risk-sensitive approach. It Models constraints on load shedding and supply-demand balancing. The paper Dall’Anese et al. (2017) directly introduces chance-constrained optimization for voltage stability and power flow Zhang and Li (2011), which is equivalent to controlling the probability of instability; CVaR and probability bounds are discussed. In 2012, Roald et al. (2014) introduces risk-constrained OPF formulations using CVaR and chance constraints. It ensures that the probability of voltage violations and instabilities is below a prescribed risk level. A year earlier in 2011, Wang et al. (2025) models the variance and tail risk of power system instability due to fluctuating wind generation. While not CVaR directly, it motivates variance and higher-order moment-based risk constraints.
@@
Multi-objective viewpoint: SafeRL can be seen as a multi-objective optimization where one objective is reward and others are (negative) costs Horie et al. (2019); Gu et al. (2025). The constraint formulation picks one point on the Pareto frontier by treating costs as hard constraints. Alternatively, one could combine reward and costs into a single scalar reward via weighted sum (penalty method), but that requires tuning weights and does not guarantee constraint satisfaction Achiam et al. (2017). Constrained formulation cleanly separates objectives and safety.